OpenStudy (anonymous):

Find f(g(x)) given: f(x) = 2x – 8 and g(x) = -5x – 10

OpenStudy (amistre64):

replace the "x" in f(x) ... with g(x)

OpenStudy (amistre64):

f(g(x)) = 2(g(x)) - 8 = 2(-5x-10) -8

OpenStudy (anonymous):

f(g(x)) = 2(-5x-10) -8 = -10x-20-8 = -10x -28